# High-Precision Measurement of D(, ) Photodisintegration Reaction and Implications for Big-Bang Nucleosynthesis

### v1

@@ Acknowledgments
We thank the staff of the Shanghai Synchrotron Radiation Facility (BL03SSID, https://cstr.cn/31124.02.SSRF.BL03SSID) for the assistance on this measurement. J. J. H. acknowledges Hiroaki Utsunomiya’s contribution to data reduction. This work was financially supported by the National Key R&D Program of China (Nos. 2022YFA1602401, 2022YFA1602400) and the National Natural Science Foundation of China (Nos. 12322509, 12275338, 12147101, 12335009, 12435010). Y. L. acknowledges the support of the Boya fellowship of Peking University and the China Postdoctoral Science Foundation (No. 2025T180924). S.-I. A. acknowledges the support of the National Research Foundation of Korea (NRF) grant (No. 2022R1F1A1070060 and 2023R1A2C1003177).

### v2

@@ Acknowledgments
We thank the staff of the Shanghai Synchrotron Radiation Facility (BL03SSID, https://cstr.cn/31124.02.SSRF.BL03SSID) for the assistance on this measurement. J. J. H. acknowledges Hiroaki Utsunomiya’s contribution to data reduction. This work was financially supported by the National Key R&D Program of China (Nos. 2022YFA1602401) and the National Natural Science Foundation of China (Nos. 12322509, 12547102, 12275338, 12335009, 12435010). Y. L. acknowledges the support of the Boya fellowship of Peking University and the China Postdoctoral Science Foundation (No. 2025T180924). S.-I. A. acknowledges the support of National Research Foundation of Korea (NRF) grant (Nos. RS-2025-16065411, 2022R1F1A1070060 and 2023R1A2C1003177).